Two trucks Truck A and truck B are driving across the country taking the same route. After 1 day truck A was 117 miles behind truck B. The next day truck A drove 185 miles more than truck B. The third day truck B drove 48 miles more miles than truck A. Where is truck A in relation to truck B after the third day?

Answer by ikleyn(53765)

.
Let's write distances:


Day 1 :  A1 = B1 - 117    (1)

Day 2 :  A2 = B2 + 185    (2)

Day 3 :  A3 = B3 -  48    (3)

-------------------------------  Add equations (1), (2) and (3).  You will get


        A1 + A2 + A3 = B1 + B2 + B3 + (-117 + 185 - 48),     or


        A1 + A2 + A3 = B1 + B2 + B3 + 20


It means that in three days Truck A covered the distance 20 miles longer than Truck B.


In other words,  Truck A is 20 miles ahead of Truck B after 3 days driving.    ANSWER

Solved.

It is a case and a good example when the Algebra language works better and is more clear than a human language.